OpenStudy (anonymous):

-2(p-4)-3p=4-3(2+p)-(-2)

OpenStudy (anonymous):

-2(p-4)-3p=4-3(2+p)-(-2) -2p+4-3p=4-6-3p+2 -5p+4=-2+2-3p -2p+4=0 4=2p p=2

OpenStudy (anonymous):

how did you get +4 in the beginning ? isn't -2x-4= *?

OpenStudy (anonymous):

It's -2*(p-4), and remember, a negative times a negative is positive!

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

Sorry, my calculations are wrong